OpenStudy (anonymous):

Find the dimension of and the basis of the kernel: L(x)=(x2, x3)^T which is a linear transformation from R^3 to R^2.

OpenStudy (anonymous):

@amistre64 do you remember anything about the above^

OpenStudy (amistre64):

the words all sound familiar. isnt a kernel related to the nulspace?

OpenStudy (anonymous):

I'm not entirely sure. I'm looking for information outside my textbook. I'm completely at a loss at this point.

OpenStudy (anonymous):

http://www.math.purdue.edu/~beranger/262/ch5/5-3.pdf

OpenStudy (anonymous):

That's what I'm looking at now.

OpenStudy (amistre64):

the dimension of the kernal, and the basis of the kernel ... so we need to determine 2 of its properties

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the transformation is a coefficient matrix = (1,1)^T ... Then the rref = (1,0)^T . So the dimension is 1 ? I think

OpenStudy (anonymous):

and the basis is (1,0)^T

OpenStudy (anonymous):

but that doesn't seem right

OpenStudy (amistre64):

ill have more time and a clearer head for this later on ... ill read up and try to refresh my memory in the mean time.

OpenStudy (anonymous):

okay!

OpenStudy (amistre64):

the kernel is related to: \[\begin{pmatrix}x^2\\x^3\end{pmatrix}\to ax^2+bx^3=0\]

OpenStudy (amistre64):

but yeah, ill chk this later ...

OpenStudy (amistre64):

http://www.math.lsa.umich.edu/~hochster/419/ker.im.html this seems about my level of comprehension today :)

OpenStudy (amistre64):

R^3 -> R^2, refers us to a 3col, 2row, it also would seem to address a setup such that L(x,y,z) , 3 elements to express R^3, ... = (x^2,x^3) , 2 elements to express R^2. http://xmlearning.maths.ed.ac.uk/lecture_notes/linear_maps/kernels_images/kernels_images_3.php L(x) = (x^2, x^3) seems to represent R -> R^2

OpenStudy (amistre64):

is that:\[a.~~L(x)=(x^2,x^3)^T\\OR\\b.~~L(x)=(x_2,x_3)^T\]

OpenStudy (amistre64):

if its polynomials (x^2, x^3), which the notation would be P->P i believe, we would be finding: 0x^3+1x^2+0x+0 = 0 1x^3+0x^2+0x+0 = 0 if its just vectors for vectors sake (x_2, x_3), then the notation R->R would make more sense to me, and thus finding: \[0x_3+1x^2+0x_1 = 0\\ 1x_3+0x_2+0x_1 = 0\]